OpenStudy (anonymous):

To buy both a new car and a new house, Tina sought two loans totaling $332,888. The simple interest rate on the first loan was 8.9%, while the simple interest rate on the second loan was 6.0%. At the end of the first year, Tina paid a combined interest payment of $20,230.51. What were the amounts of the two loans?

OpenStudy (anonymous):

\[A+B=332888,1.089A+1.06B=20230.51\]is how you express the problem mathematically. I think you can do it now?

OpenStudy (anonymous):

Oh, I'm sorry, since Tina's paying back ONLY the interest, and not the entire loan,\[A+B=332888,0.089A+0.06B=20230.51\]